# ShieldAgent: Shielding Agents via Verifiable Safety Policy Reasoning

### v1

@@ Abstract
Autonomous agents powered by foundation models have seen widespread adoption across various real-world applications. However, they remain highly vulnerable to malicious instructions and attacks, which can result in severe consequences such as privacy breaches and financial losses. More critically, existing guardrails for LLMs are not applicable due to the complex and dynamic nature of agents. To tackle these challenges, we propose ShieldAgent, the first guardrail agent designed to enforce explicit safety policy compliance for the action trajectory of other protected agents through logical reasoning. Specifically, ShieldAgent first constructs a safety policy model by extracting verifiable rules from policy documents and structuring them into a set of action-based probabilistic rule circuits. Given the action trajectory of the protected agent, ShieldAgent retrieves relevant rule circuits and generates a shielding plan, leveraging its comprehensive tool library and executable code for formal verification. In addition, given the lack of guardrail benchmarks for agents, we introduce ShieldAgent-Bench, a dataset with 3K safety-related pairs of agent instructions and action trajectories, collected via SOTA attacks across 6 web environments and 7 risk categories. Experiments show that ShieldAgent achieves SOTA on ShieldAgent-Bench and three existing benchmarks, outperforming prior methods by $11.3\%$ on average with a high recall of $90.1\%$. Additionally, ShieldAgent reduces API queries by $64.7\%$ and inference time by $58.2\%$, demonstrating its high precision and efficiency in safeguarding agents.
@@ 1 Introduction
LLM-based autonomous agents are rapidly gathering momentum across various applications, integrating their ability to call external tools and make autonomous decisions in real-world tasks such as web browsing (Zhou et al., 2023), GUI navigation (Lin et al., 2024), and embodied control (Mao et al., 2023). Among these, LLM-based web agents, such as OpenAI’s Operator (OpenAI, 2025b), deep research agent (OpenAI, 2025a), and Anthropic’s computer assistant agent (Anthropic, 2024), have become particularly prominent, driving automation in areas like online shopping, stock trading, and information retrieval.
@@
Despite their growing capabilities, users remain reluctant to trust current web agents with high-stakes data and assets, as they are still highly vulnerable to malicious instructions and adversarial attacks (Chen et al., 2024c; Wu et al., 2025), which can lead to severe consequences such as privacy breaches and financial losses (Levy et al., 2024). Existing guardrails primarily focus on LLMs as models, while failing to safeguard them as agentic systems due to two key challenges: (1) LLM-based agents operate through sequential interactions with dynamic environments, making it difficult to capture unsafe behaviors that emerge over time (Xiang et al., 2024); (2) Safety policies governing these agents are often complex and encoded in lengthy regulation documents (e.g. EU AI Act (Act, 2024)) or corporate policy handbooks (GitLab, 2025), making it difficult to systematically extract, verify, and enforce rules across different platforms (Zeng et al., 2024). As a result, safeguarding the safety of LLM-based web agents remains an open challenge.
@@
To address these challenges, we introduce ShieldAgent, the first LLM-based guardrail agent designed to shield the action trajectories of other LLM-based autonomous agents, ensuring explicit safety compliance through probabilistic logic reasoning and verification. Unlike existing approaches that rely on simple text-based filtering (Xiang et al., 2024), ShieldAgent accounts for the uniqueness of agent actions and explicitly verifies them against relevant policies in an efficient manner. At its core, ShieldAgent automatically constructs a robust safety policy model by extracting verifiable rules from policy documents, iteratively refining them, and grouping them based on different action types to form a set of structured, action-based probabilistic rule circuits (Kang & Li, 2024). During inference, ShieldAgent only verifies the relevant rule circuits corresponding to the invoked action, ensuring both precision and efficiency. Specifically, ShieldAgent references from a hybrid memory module of both long-term shielding workflows and short-term interaction history, generates a shielding plan with specialized operations from a rich tool library, and runs formal verification code. Once a rule is verified, ShieldAgent performs probabilistic inference within the circuits and provides a binary safety label, identifies any violated rules, and generates detailed explanations to justify its decision.
@@
While evaluating these guardrails is critical for ensuring agent safety, existing benchmarks remain small in scale, cover limited risk categories, and lack explicit risk definitions (see Table 1). Therefore, we introduce ShieldAgent-Bench, the first comprehensive agent guardrail benchmark comprising 2K safety-related pairs of agent instructions and trajectories across six web environments and seven risk categories. Specifically, each unsafe agent trajectory is generated under two types of attacks (Chen et al., 2024c; Xu et al., 2024) based on different perturbation sources (i.e., agent-based and environment-based), capturing risks present both within the agent system and the external environments.
@@
We conduct extensive experiments demonstrating that ShieldAgent achieves SOTA performance on both ShieldAgent-Bench and three existing benchmarks (i.e., ST-WebAgentBench (Levy et al., 2024), VWA-Adv (Wu et al., 2025), and AgentHarm (Andriushchenko et al., )). Specifically, ShieldAgent outperforms the previous best guardrail method by $11.3\%$ on ShieldAgent-Bench, and $7.4\%$ on average across existing benchmarks. Grounded on robust safety policy reasoning, it achieves the lowest false positive rate at $4.8\%$ and a high recall rate of violated rules at $90.1\%$. Additionally, ShieldAgent reduces the number of closed-source API queries by $64.7\%$ and inference time by $58.2\%$, demonstrating its ability to effectively shield LLM agents’ actions while significantly improving efficiency and reducing computational overhead.
@@ 2.1 Safety of LLM Agents
While LLM agents are becoming increasingly capable, numerous studies have demonstrated their susceptibility to manipulated instructions and vulnerability to adversarial attacks, which often result in unsafe or malicious actions (Levy et al., 2024; Andriushchenko et al., ; Zhang et al., 2024b). Existing attack strategies against LLM agents can be broadly classified into the following two categories.
(1) Agent-based attacks, where adversaries manipulate internal components of the agent, such as instructions (Guo et al., ; Zhang et al., 2024c), memory modules or knowledge bases (Chen et al., 2024c; Jiang et al., 2024), and tool libraries (Fu et al., 2024; Zhang et al., 2024a). These attacks are highly effective and can force the agent to execute arbitrary malicious requests. However, they typically require some access to the agent’s internal systems or training data.
(2) Environment-based attacks, which exploit vulnerabilities in the environment that the agents interact with to manipulate their behavior (Liao et al., 2024), such as injecting malicious HTML elements (Xu et al., 2024) or deceptive web pop-ups (Zhang et al., 2024d). Since the environment is less controlled than the agent itself, these attacks are easier to execute in real world but may have a lower success rate.
Both attack types pose significant risks, leading to severe consequences such as life-threatening failures (Chen et al., 2024c), privacy breaches (Liao et al., 2024), and financial losses (Andriushchenko et al., ). Therefore in this work, we account for both agent-based and environment-based adversarial perturbations in the design of ShieldAgent. Besides, we leverage SOTA attacks (Chen et al., 2024c; Xu et al., 2024) from both categories to construct our ShieldAgent-Bench dataset which involves diverse risky web agent trajectories across various environments.
@@ 2.2 LLM Guardrails
While LLM agents are highly vulnerable to adversarial attacks, existing guardrail mechanisms are designed for LLMs as models rather than agents, leaving a critical gap in safeguarding their sequential decision-making processes (Andriushchenko et al., ). Current guardrails primarily focus on filtering harmful inputs and outputs, such as LlamaGuard (Inan et al., 2023) for text-based LLMs, LlavaGuard (Helff et al., 2024) for image-based multimodal LLMs, and SafeWatch (Chen et al., 2024a) for video generative models. However, these methods focus solely on content moderation, failing to address the complexities of action sequences, where vulnerabilities often emerge over time (Debenedetti et al., 2024). While GuardAgent (Xiang et al., 2024) preliminarily explores the challenge of guardrailing LLM agents with another LLM agent, it focus solely on textual space and still relies on the model’s internal knowledge rather than explicitly enforcing compliance with external safety policies and regulations (Zeng et al., 2024), limiting its effectiveness in real-world applications. To our knowledge, ShieldAgent is the first multimodal LLM-based agent to safeguard action sequences of other LLM agents via probabilistic policy reasoning to ensure explicit and efficient policy compliance.
@@ 3.2.1 Ovewview of ASPM
Specifically, all constraints are represented as linear temporal logic (LTL) rules (Zhu et al., 2017) where each rule includes corresponding atomic predicates as decision variables [^1] Each predicate can be assigned a boolean value per time step to describe the agent system variables or environment state.. Please refer to §3.2.2 for details. Thus let $\mathcal{P}$, $\mathcal{R}$ denote the predicate and rule space respectively, we have:
@@
where $\pi_{\theta}$ denotes the probabilistic logic model (parameterized by $\theta$) which organizes the rules (see §3.2.4). Specifically, $\mathcal{G}_{\text{ASPM}}$ partitions $\mathcal{P}$ into state predicates $p_{\text{s}}\in\mathcal{P}_{\text{s}}$ to represent system states or environmental conditions, and action predicates $p_{\text{a}}\in\mathcal{P}_{\text{a}}$ to represent target actions. Consequently, $\mathcal{R}$ is divided into action rules $\mathcal{R}_{\text{a}}$ which encodes safety specifications for target actions, and physical rules $\mathcal{R}_{\text{p}}$ which capture internal constraints on system variables. Specifically, while $\mathcal{R}_{\text{p}}$ does not directly constrain actions in $\mathcal{P}_{\text{a}}$, these knowledge rules are critical for the logical reasoning in ASPM, enhancing the robustness of our shield (Kang & Li, 2024). Therefore, by structuring the solution space this way, we achieve a clear and manageable verification of target actions. Refer to Appendix A.2 for more details.
Specifically, we construct ASPM from policy documents via the following steps: (1) Extract structured safety rules from government regulations (Act, 2024), corporate policies (GitLab, 2025), and user-provided constraints; (2) Refine these rules iteratively for better clarity, verifiability, and efficiency; (3) Cluster the optimized rules by different agent actions and obtain a set of action-based rule circuits (Kisa et al., 2014) where each circuit associates an agent action with relevant rules for verification; (4) Train the ASPM by learning rule weights from either real-world interactions or simulated data, ensuring adaptive and robust policy verification.
@@ 3.2.2 Automatic Policy and Rule Extraction
Since policy definitions are typically encoded in lengthy documents with structures varying widely across platforms (Act, 2024; GitLab, 2025), directly verifying them is challenging. To address this, ShieldAgent first extracts individual actionable policies from these documents and further translates them into manageable logical rules for tractable verification.
@@ 3.2.3 ASPM Structure Optimization
$\mathcal{V}_{i}=\frac{1}{k}\sum_{i=m}^{k}S_{\alpha(m)}\>\>\text{s.t.}\>\>\,S_{\alpha}=\mathrm{desc}\bigl{(}\{e_{i}\cdot e_{j}\mid j\leq|\mathcal{P}_{i}|\}\bigr{)}\vspace{-0.05in}$ (3)
where $e_{i}$ denotes the normalized vector representation of predicate $p_{i}$ obtained by a SOTA embedding model (e.g. OpenAI’s text-embedding-3-large model (OpenAI, 2024)). Please refer to Appendix C.4 for more details.
@@ 3.2.4 ASPM Inference & Training
Action-based ASPM Clustering. Observing that certain agent actions exhibit low logical correlation to each other (e.g. $delete\_data$ and $buy\_product$), we further construct an action-based probabilistic circuit $\pi_{\theta}$ (Kisa et al., 2014) from ASPM to boost its verification efficiency while retaining precision. Concretely, we first apply spectral clustering (Von Luxburg, 2007) to the state predicates $\mathcal{P}_{s}$, grouping rules that exhibit strong logical dependencies or high semantic relevance. Then, we associate each action predicate $p_{a}$ with its relevant constraints by unifying rule clusters that involve $p_{a}$ into a single probabilistic circuit $\mathcal{C}_{\theta_{a}}^{p_{a}}$ (weights $\theta_{a}$ are trained in §3.2.4). During verification, the agent only needs to check the corresponding circuit w.r.t. the invoked action, thereby substantially reducing inference complexity while preserving logical dependencies among rules.
@@
In each action circuit $\mathcal{C}_{\theta_{a}}^{p_{a}}$, the joint distribution over all possible assignments of predicates (i.e., world) is modeled via Markov Logic Network (Richardson & Domingos, 2006). Let $\mu_{p}$ denote the assignment of predicate $p$, the probability of the proposed world $\mu$ with action $p_{a}$ invoked is given by:
@@
where $\mathbb{I}[\mu\sim r]=1$ indicates that the world $\mu$ follows the logical rule $r$ and $Z$ is a constant partition for normalization. However, since the absolute value of world probability is usually unstable (Gürel et al., 2021), directly thresholding it as the guardrail decision may cause a high false positive rate. Thus inspired by the control barrier certificate (Ames et al., 2019), we propose the following relative safety condition:
@@
where labels $y^{(i)}=1$ if action $a^{(i)}$ is safe or $y^{(i)}=-1$ if unsafe. Specifically, $y^{(i)}$ can be derived from either real-world safety-labeled data or simulated pseudo-learning (Kang & Li, 2024). The learned weights act as soft constraints, capturing the relative importance of each rule in guardrail decision-making. We illustrate the training process in Algorithm 3.
@@
1:  $p_{a}\leftarrow\textsc{Extract}(a_{i})$triangleright Extract action predicates
2:  $\mathcal{C}_{\theta_{a}}^{p_{a}}=\bigl{(}\mathcal{P}_{p_{a}},R_{p_{a}},\theta_{a}\bigr{)}\leftarrow\textsc{Retrieve}(p_{a},\,\mathcal{G}_{\text{ASPM}})$
3:  $\mathcal{V}_{s}=\{p_{s}^{i}:v_{s}^{i}\}\leftarrow\emptyset$triangleright Initialize predicate-value map
@@
7:        $A_{s}\leftarrow\textsc{Plan}(\mathcal{W}_{r},r,\mathcal{P}_{r})$triangleright Generate an action plan with shielding operations (e.g., Search, Check)
@@
9:           $o_{s}^{i}\leftarrow\textsc{Execute}(t_{s}^{i},\mathcal{H}_{<i},o_{i})$triangleright Get step result
10:           $\mathcal{V}_{s}[p_{s}]\leftarrow\textsc{Parse}\bigl{(}o_{s}^{i}\bigr{)},p_{s}\in\mathcal{P}_{r}$triangleright Attempt to assign a truth value to any unassigned predicates
@@
15:  $\epsilon_{s}\leftarrow P_{\theta}\bigl{(}\mu_{p_{a}=1}\bigr{)}-P_{\theta}\bigl{(}\mu_{p_{a}=0}\bigr{)}$triangleright Calculate safety condition via Eq. (4) and Eq. (5)
@@
17:     $l_{s}\leftarrow 1$triangleright Action $p_{a}$ is safe
@@
19:     $l_{s}\leftarrow 0$triangleright Action $p_{a}$ is unsafe
@@
21:  return $\bigl{(}l_{s},V_{s},T_{s}\bigr{)}$triangleright Return safety label, violated rules, textual explanation
@@
ST-Web (Levy et al., 2024) 234 3 0 3 36
AgentHarm (Andriushchenko et al., ) 440 1 0 0 0
VWA-Adv (Wu et al., 2025) 200 1 1 3 0

### v2

@@ Abstract
Autonomous agents powered by foundation models have seen widespread adoption across various real-world applications. However, they remain highly vulnerable to malicious instructions and attacks, which can result in severe consequences such as privacy breaches and financial losses. More critically, existing guardrails for LLMs are not applicable due to the complex and dynamic nature of agents. To tackle these challenges, we propose ShieldAgent, the first guardrail agent designed to enforce explicit safety policy compliance for the action trajectory of other protected agents through logical reasoning. Specifically, ShieldAgent first constructs a safety policy model by extracting verifiable rules from policy documents and structuring them into a set of action-based probabilistic rule circuits. Given the action trajectory of the protected agent, ShieldAgent retrieves relevant rule circuits and generates a shielding plan, leveraging its comprehensive tool library and executable code for formal verification. In addition, given the lack of guardrail benchmarks for agents, we introduce ShieldAgent-Bench, a dataset with 3K safety-related pairs of agent instructions and action trajectories, collected via SOTA attacks across 6 web environments and 7 risk categories. Experiments show that ShieldAgent achieves SOTA on ShieldAgent-Bench and three existing benchmarks, outperforming prior methods by $11.3\%$ on average with a high recall of $90.1\%$. Additionally, ShieldAgent reduces API queries by $64.7\%$ and inference time by $58.2\%$, demonstrating its high precision and efficiency in safeguarding agents. Our project is available and continuously maintained here: https://shieldagent-aiguard.github.io/
@@ 1 Introduction
LLM-based autonomous agents are rapidly gathering momentum across various applications, integrating their ability to call external tools and make autonomous decisions in real-world tasks such as web browsing (zhou2023webarena), GUI navigation (lin2024showui), and embodied control (mao2023language). Among these, LLM-based web agents, such as OpenAI’s Operator (operator2025), deep research agent (deepresearch2025), and Anthropic’s computer assistant agent (mcp2025), have become particularly prominent, driving automation in areas like online shopping, stock trading, and information retrieval.
@@
Despite their growing capabilities, users remain reluctant to trust current web agents with high-stakes data and assets, as they are still highly vulnerable to malicious instructions and adversarial attacks (chenagentpoison; wudissecting), which can lead to severe consequences such as privacy breaches and financial losses (levy2024st). Existing guardrails primarily focus on LLMs as models, while failing to safeguard them as agentic systems due to two key challenges: (1) LLM-based agents operate through sequential interactions with dynamic environments, making it difficult to capture unsafe behaviors that emerge over time (xiang2024guardagent); (2) Safety policies governing these agents are often complex and encoded in lengthy regulation documents (e.g. EU AI Act (act2024eu)) or corporate policy handbooks (gitlab2025), making it difficult to systematically extract, verify, and enforce rules across different platforms (zeng2024air). As a result, safeguarding the safety of LLM-based web agents remains an open challenge.
@@
To address these challenges, we introduce ShieldAgent, the first LLM-based guardrail agent designed to shield the action trajectories of other LLM-based autonomous agents, ensuring explicit safety compliance through probabilistic logic reasoning and verification. Unlike existing approaches that rely on simple text-based filtering (xiang2024guardagent), ShieldAgent accounts for the uniqueness of agent actions and explicitly verifies them against relevant policies in an efficient manner. At its core, ShieldAgent automatically constructs a robust safety policy model by extracting verifiable rules from policy documents, iteratively refining them, and grouping them based on different action types to form a set of structured, action-based probabilistic rule circuits (kang2024r). During inference, ShieldAgent only verifies the relevant rule circuits corresponding to the invoked action, ensuring both precision and efficiency. Specifically, ShieldAgent references from a hybrid memory module of both long-term shielding workflows and short-term interaction history, generates a shielding plan with specialized operations from a rich tool library, and runs formal verification code. Once a rule is verified, ShieldAgent performs probabilistic inference within the circuits and provides a binary safety label, identifies any violated rules, and generates detailed explanations to justify its decision.
@@
While evaluating these guardrails is critical for ensuring agent safety, existing benchmarks remain small in scale, cover limited risk categories, and lack explicit risk definitions (see Table 1). Therefore, we introduce ShieldAgent-Bench, the first comprehensive agent guardrail benchmark comprising 2K safety-related pairs of agent instructions and trajectories across six web environments and seven risk categories. Specifically, each unsafe agent trajectory is generated under two types of attacks (chenagentpoison; xu2024advweb) based on different perturbation sources (i.e., agent-based and environment-based), capturing risks present both within the agent system and the external environments.
@@
We conduct extensive experiments demonstrating that ShieldAgent achieves SOTA performance on both ShieldAgent-Bench and three existing benchmarks (i.e., ST-WebAgentBench (levy2024st), VWA-Adv (wudissecting), and AgentHarm (andriushchenkoagentharm)). Specifically, ShieldAgent outperforms the previous best guardrail method by $11.3\%$ on ShieldAgent-Bench, and $7.4\%$ on average across existing benchmarks. Grounded on robust safety policy reasoning, it achieves the lowest false positive rate at $4.8\%$ and a high recall rate of violated rules at $90.1\%$. Additionally, ShieldAgent reduces the number of closed-source API queries by $64.7\%$ and inference time by $58.2\%$, demonstrating its ability to effectively shield LLM agents’ actions while significantly improving efficiency and reducing computational overhead.
@@ 2.1 Safety of LLM Agents
While LLM agents are becoming increasingly capable, numerous studies have demonstrated their susceptibility to manipulated instructions and vulnerability to adversarial attacks, which often result in unsafe or malicious actions (levy2024st; andriushchenkoagentharm; zhang2024agent). Existing attack strategies against LLM agents can be broadly classified into the following two categories.
(1) Agent-based attacks, where adversaries manipulate internal components of the agent, such as instructions (guoredcode; zhang2024towards), memory modules or knowledge bases (chenagentpoison; jiang2024rag), and tool libraries (fu2024imprompter; zhang2024breaking). These attacks are highly effective and can force the agent to execute arbitrary malicious requests. However, they typically require some access to the agent’s internal systems or training data.
(2) Environment-based attacks, which exploit vulnerabilities in the environment that the agents interact with to manipulate their behavior (liao2024eia), such as injecting malicious HTML elements (xu2024advweb) or deceptive web pop-ups (zhang2024attacking). Since the environment is less controlled than the agent itself, these attacks are easier to execute in real world but may have a lower success rate.
Both attack types pose significant risks, leading to severe consequences such as life-threatening failures (chenagentpoison), privacy breaches (liao2024eia), and financial losses (andriushchenkoagentharm). Therefore in this work, we account for both agent-based and environment-based adversarial perturbations in the design of ShieldAgent. Besides, we leverage SOTA attacks (chenagentpoison; xu2024advweb) from both categories to construct our ShieldAgent-Bench dataset which involves diverse risky web agent trajectories across various environments.
@@ 2.2 LLM Guardrails
While LLM agents are highly vulnerable to adversarial attacks, existing guardrail mechanisms are designed for LLMs as models rather than agents, leaving a critical gap in safeguarding their sequential decision-making processes (andriushchenkoagentharm). Current guardrails primarily focus on filtering harmful inputs and outputs, such as LlamaGuard (inan2023llama) for text-based LLMs, LlavaGuard (helff2024llavaguard) for image-based multimodal LLMs, and SafeWatch (chen2024safewatch) for video generative models. However, these methods focus solely on content moderation, failing to address the complexities of action sequences, where vulnerabilities often emerge over time (debenedetti2024agentdojo). While GuardAgent (xiang2024guardagent) preliminarily explores the challenge of guardrailing LLM agents with another LLM agent, it focus solely on textual space and still relies on the model’s internal knowledge rather than explicitly enforcing compliance with external safety policies and regulations (zeng2024air), limiting its effectiveness in real-world applications. To our knowledge, ShieldAgent is the first multimodal LLM-based agent to safeguard action sequences of other LLM agents via probabilistic policy reasoning to ensure explicit and efficient policy compliance.
@@ 3.2.1 Ovewview of ASPM
Specifically, all constraints are represented as linear temporal logic (LTL) rules (zhu2017symbolic) where each rule includes corresponding atomic predicates as decision variables [^1] Each predicate can be assigned a boolean value per time step to describe the agent system variables or environment state.. Please refer to §3.2.2 for details. Thus let $\mathcal{P}$, $\mathcal{R}$ denote the predicate and rule space respectively, we have:
@@
where $\pi_{\theta}$ denotes the probabilistic logic model (parameterized by $\theta$) which organizes the rules (see §3.2.4). Specifically, $\mathcal{G}_{\text{ASPM}}$ partitions $\mathcal{P}$ into state predicates $p_{\text{s}}\in\mathcal{P}_{\text{s}}$ to represent system states or environmental conditions, and action predicates $p_{\text{a}}\in\mathcal{P}_{\text{a}}$ to represent target actions. Consequently, $\mathcal{R}$ is divided into action rules $\mathcal{R}_{\text{a}}$ which encodes safety specifications for target actions, and physical rules $\mathcal{R}_{\text{p}}$ which capture internal constraints on system variables. Specifically, while $\mathcal{R}_{\text{p}}$ does not directly constrain actions in $\mathcal{P}_{\text{a}}$, these knowledge rules are critical for the logical reasoning in ASPM, enhancing the robustness of our shield (kang2024r). Therefore, by structuring the solution space this way, we achieve a clear and manageable verification of target actions. Refer to Appendix A.2 for more details.
Specifically, we construct ASPM from policy documents via the following steps: (1) Extract structured safety rules from government regulations (act2024eu), corporate policies (gitlab2025), and user-provided constraints; (2) Refine these rules iteratively for better clarity, verifiability, and efficiency; (3) Cluster the optimized rules by different agent actions and obtain a set of action-based rule circuits (kisa2014probabilistic) where each circuit associates an agent action with relevant rules for verification; (4) Train the ASPM by learning rule weights from either real-world interactions or simulated data, ensuring adaptive and robust policy verification.
@@ 3.2.2 Automatic Policy and Rule Extraction
Since policy definitions are typically encoded in lengthy documents with structures varying widely across platforms (act2024eu; gitlab2025), directly verifying them is challenging. To address this, ShieldAgent first extracts individual actionable policies from these documents and further translates them into manageable logical rules for tractable verification.
@@ 3.2.3 ASPM Structure Optimization
$\mathcal{V}_{i}=\frac{1}{k}\sum_{i=m}^{k}S_{\alpha(m)}\>\>\text{s.t.}\>\>\,S_{\alpha}=\mathrm{desc}\bigl(\{e_{i}\cdot e_{j}\mid j\leq|\mathcal{P}_{i}|\}\bigr)\vskip-3.61371pt$ (3)
where $e_{i}$ denotes the normalized vector representation of predicate $p_{i}$ obtained by a SOTA embedding model (e.g. OpenAI’s text-embedding-3-large model (openaiembedding2024)). Please refer to Appendix C.4 for more details.
@@ 3.2.4 ASPM Inference & Training
Action-based ASPM Clustering. Observing that certain agent actions exhibit low logical correlation to each other (e.g. delete_datadelete\_data and buy_productbuy\_product), we further construct an action-based probabilistic circuit $\pi_{\theta}$ (kisa2014probabilistic) from ASPM to boost its verification efficiency while retaining precision. Concretely, we first apply spectral clustering (von2007tutorial) to the state predicates $\mathcal{P}_{s}$, grouping rules that exhibit strong logical dependencies or high semantic relevance. Then, we associate each action predicate $p_{a}$ with its relevant constraints by unifying rule clusters that involve $p_{a}$ into a single probabilistic circuit $\mathcal{C}_{\theta_{a}}^{p_{a}}$ (weights $\theta_{a}$ are trained in §3.2.4). During verification, the agent only needs to check the corresponding circuit w.r.t. the invoked action, thereby substantially reducing inference complexity while preserving logical dependencies among rules.
@@
In each action circuit $\mathcal{C}_{\theta_{a}}^{p_{a}}$, the joint distribution over all possible assignments of predicates (i.e., world) is modeled via Markov Logic Network (richardson2006markov). Let $\mu_{p}$ denote the assignment of predicate $p$, the probability of the proposed world $\mu$ with action $p_{a}$ invoked is given by:
@@
where $\mathbb{I}[\mu\sim r]=1$ indicates that the world $\mu$ follows the logical rule $r$ and $Z$ is a constant partition for normalization. However, since the absolute value of world probability is usually unstable (gurel2021knowledge), directly thresholding it as the guardrail decision may cause a high false positive rate. Thus inspired by the control barrier certificate (ames2019control), we propose the following relative safety condition:
@@
where labels $y^{(i)}=1$ if action $a^{(i)}$ is safe or $y^{(i)}=-1$ if unsafe. Specifically, $y^{(i)}$ can be derived from either real-world safety-labeled data or simulated pseudo-learning (kang2024r). The learned weights act as soft constraints, capturing the relative importance of each rule in guardrail decision-making. We illustrate the training process in Algorithm 3.
@@
[…]
@@
7:   $A_{s}\leftarrow\textsc{Plan}(\mathcal{W}_{r},r,\mathcal{P}_{r})$ $\triangleright$ Generate an action plan with shielding operations (e.g., Search, Check)
@@
9:    $o_{s}^{i}\leftarrow\textsc{Execute}(t_{s}^{i},\mathcal{H}_{<i},o_{i})$ $\triangleright$ Get step result
10:    $\mathcal{V}_{s}[p_{s}]\leftarrow\textsc{Parse}\bigl(o_{s}^{i}\bigr),p_{s}\in\mathcal{P}_{r}$ $\triangleright$ Attempt to assign a truth value to any unassigned predicates
@@
[…]
@@
17:  $l_{s}\leftarrow 1$ $\triangleright$ Action $p_{a}$ is safe
@@
19:  $l_{s}\leftarrow 0$ $\triangleright$ Action $p_{a}$ is unsafe
@@
21:return $\bigl(l_{s},V_{s},T_{s}\bigr)$ $\triangleright$ Return safety label, violated rules, textual explanation
@@
ST-Web (levy2024st) 234 3 0 3 36
AgentHarm (andriushchenkoagentharm) 440 1 0 0 0
VWA-Adv (wudissecting) 200 1 1 3 0